OpenStudy (anonymous):

College Algebra: reviewing for Final Exam!! Which one of the following does not define y as a function of x? A.) x^2+y=16 B.) y=(x-3)^2-1 C.) x+y=3 D.) x=y^2

OpenStudy (anonymous):

@dan815 @freckles @jim_thompson5910 @Nnesha @TheSmartOne @triciaal Any of you able to help me? :)

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

No where. I don't even know where to start.

OpenStudy (triciaal):

y as a function of x means you have y = whatever

OpenStudy (anonymous):

Oh. Is it always y=whatever?

jimthompson5910 (jim_thompson5910):

if you solved one of those equations for y, you'll get y = plus/minus... that will be the equation that is NOT a function because y = the plus form while it also equals the negative form. That means one x value could produce 2 y values

jimthompson5910 (jim_thompson5910):

Rule: a relation is a function if every given x input produces EXACTLY ONE output y.

OpenStudy (anonymous):

Hmmm

OpenStudy (triciaal):

@Natasha.g.2013 @jim_thompson5910 last post is the formal way to say it

OpenStudy (triciaal):

i already gave him the medal

OpenStudy (anonymous):

He's right, I'm not saying he isn't. I was writing down the rule, so I wouldn't forget it.

OpenStudy (anonymous):

So, the answer would be D.) x=y^2 because it can only produce 1 y

OpenStudy (triciaal):

he is a very reliable source

OpenStudy (anonymous):

Thank you very much @jim_thompson5910 and @triciaal for helping me! :)

jimthompson5910 (jim_thompson5910):

the answer is x = y^2 because solving for y gives you y = plus/minus sqrt(x) so y is either y = sqrt(x) or it is y = -sqrt(x)

OpenStudy (anonymous):

Woohoo!! :)

jimthompson5910 (jim_thompson5910):

if x = 4, then y is either +2 or -2 so that is one example showing how it's not a function

OpenStudy (anonymous):

You're right! :)

OpenStudy (triciaal):

go and ace that test !

OpenStudy (anonymous):

Thank you!!!! :)